OpenStudy (anonymous):

The governor of state A earns $51,660 more than governor of state B. If the total of their salaries is $308,500. Find the salaries of each

OpenStudy (anonymous):

Well, first thing to do is to make an equation. The salary of Governor A can be represented as n. The salary of Governor B can be represented as 51,660 + n, because he makes 51, 660 more than Governor A. We know that these two expressions added together equals 308, 500. n + n + 51660 = 308500 When you simplify this: 2n + 51660 = 308500 Now, solve for n: 2n = 308500 - 51660 2n = 256840 Divide both sides by 2: n = 128420. So the salary of Governor A (which was represented by n) is $128,420. We know that the salary of Governor B was $51,660 MORE than Governor A so his salary can be represented by 51,660 + 128420 = 180,080 Therefore, Governor A has a salary of $128,420 while Governor B has a salary of $180,080. Hope this helped!